the area of a rectangular playground is 160 sq.m. and its perimeter is 52sq.m. what is the length and width of the playground?
please help...please also show the solution so i will know how to solve...thanks

Answer by RAY100(1637)

Area = Length * Width,,A=LW = 160,,,,,,,or L=160/W
.
Perimeter = 2(L+W) ,,,,,P=2L +2W =52,,,,subst from above
.
52= 2(160/W) +2W,,,,,multiply thru by "W"
.
52W = 320 +2W^2
.
2w^2 -52W +320 =0,,,,,divide by 2
.
W^2 -26W +160 =0
.
solve for W,,,using Quadratic Formula
.
a=1,,,,b=(-26),,,c=160
.
W=[ -(-26) +/-sqrt{ (-26)^2 -4 (1)(160)}]/2(1)
.
W= [26 +/- sqrt{676 -640}]/2
.
W= [26 +/- 6]/2
.
w= 10,,16
,
solving for L=160/W,,,L= 16,10
.
Therefore,,,W=10,,,L=16,,,,,,by convention W less than L
.
Check, A= 10*16=160,,,ok
.
P=2(10+16)=2*26 = 52,,,,,ok
.